# POV Learning: Individual Alignment of Multimodal Models using Human Perception

### v1

@@ 4.2 Perception-Guided LSTM Model
For this setup, we implement a unidirectional LSTM model that solely receives symbolic representations of the AOIs as a sequence. In our case, we refer to all annotated regions from the eye tracking study as AOIs. If we take the sequence from Table 1 as an example, each row corresponds to one step in the LSTM. In this example, the symbol denoted vis_wall would correspond to the input in the first step t1, txt_wall to t2, and so on until the last element of the sequence vis_wall is reached [^4] Note that we do not add any additional information to the sequence, so that representations are purely symbolic. Before feeding the sequence into the LSTM, we put the one-hot encoded elements through an embedding layer that outputs a dense low-dimensional vector representation of the elements in the AOI sequence. Additionally, we apply the same technique to create embeddings for the participants. As a final step we concatenate each AOI sequence element with the current participant embedding and put it through a feed-forward layer to receive a combined representation, capturing Who is looking at which AOI. This final representation is then used as the input for the LSTM, of which we take the final hidden state and perform the classification by running it through a final feed-forward layer which serves as classification layer. During training, the Embedding layers are also updated, so that the resulting representations are fitted to the task. Overall, there are 837 unique AOI identifiers (i.e. the text representations like ”vis_wall” or ”txt_zebra”) in the dataset. The symbolic AOI representations are not stimulus specific, but rather on a conceptual level, which is why the ”vocabulary” is relatively small. This setting can be seen as using only the individual alignment signal, without the stimulus itself.
@@ 4.3 Content-based Transformer Model
The fundamental building block for this baseline is a transformer encoder, which comprises multiple attention heads and layers. Figure 4 shows the baseline model, which can be extended and adapted to different tasks by adding additional layers on top. We use pre-trained BERT [^5] https://huggingface.co/docs/transformers/v4.28.1/en/model_doc/bert [4] embeddings that are obtained from the Huggingface Transformers Package [27] as representations for the linguistic inputs (the captions) and a pre-trained ResNet50666https://pytorch.org/vision/main/models/resnet.html [6] for obtaining image features of the pre-defined regions. The reason for using pre-defined regions instead of applying an object detector as in most state-of-the-art approaches is to ensure a one-to-one mapping between eye tracking data and machine learning input.
@@
For the experiments described below, we additionally concatenated a low-dimensional participant representation to the cross-modal classification token. The main reason for this is to function as rudimentary alignment signal. There is also a single feed-forward layer between the ResNet image features and the transformer encoder, which is used for downsizing the image feature dimension to match with the BERT Embedding dimension. For the final decision, the classification token is put through a single feed-forward layer with three output dimensions that correspond to the three possible outcomes. Note that the ResNet and BERT weights are not updated during training. Conceptually, the main advantage of the Transformer architecture is that it can use background knowledge in form of the pre-trained representations of the input modalities to solve the Perception-guided Crossmodal Entailment task.
@@ 4.4 Ensemble Model
In addition to the perception-guided approach of the LSTM and the content-based approach of the Transformer model, we introduce a meta-model as a combination of both. In this Ensemble model, we simply combine both of the models described above and use the average of both classifications as a final classification. This has been shown sufficient in our case, where we train the Ensemble parts in a joint fashion and do not require to introduce additional trainable parameters. By training both the LSTM and Transformer model together, we assume that the model learns to differentiate between the cases where one model performs better than the other, thus it decides which of the models returns the better classification in a given case.
5 Experiments I - POV Models based on Fixation Sequences
@@
In a first set of experiments, we investigated how the different kinds of information (content vs. sequence) contribute to solving the PCE task on an individual level. For a fair comparison, all three models are trained using the same hyperparameters that are based on the best-performing Ensemble model. We use 6 attention heads and layers for the transformer, 256 dimensions for the transformer feed-forward network and 32 embedding dimensions (for both the participant and AOI embedding) in the LSTM network. All three models are trained using a batch size of 128 and a learning rate of 0.0001. We report the prediction accuracy and the F1 score (averaged over all classes) on the test split as evaluation metrics for all conducted experiments. Since the rare unclear class is problematic for purely sequential models, we also report Accuracy and F1 scores that excludes samples with this class (’2 Class’ setting). In addition to our models, we report a naive baseline that only predicts the most frequent class. Table 3 shows that the combination of sequential and contextual information in an ensemble leads to the best results.
Table 3: Experimental Results for the PCE task.
Setting Metric LSTM Transformer Ensemble Naive
3 Class Accuracy .7466 .7238 .7728 .6715
F1 .4849 .5240 .5628 .2678
2 Class Accuracy .8494 .8611 .8742 .7639
F1 .7767 .7879 .8112 .4330
@@
These results clearly confirm our hypothesis that the fixation sequence contains information that can be used as alignment signal in a POV learning task. Additionally, they show that the combination of background and sequential information in an Ensemble model leads to the best results.
To test the importance of providing a participant representation to aid the alignment we investigate how much the individual participant embeddings (see Sec. 4.2) of our models contribute to their overall performance. We want to find out whether it is sufficient to use individual behavior (the fixations sequences) for alignment, or if the models are improved by a dedicated representation of an individual. For this experiment, we use the same hyperparameters that were used to achieve the results in table 3. The only difference is that here we removed all individual participant’s representations from the models. We report the same evaluation metrics as before and show the relative change in percent.
Table 4: Ablation Results for the PCE task without Participant Embedding.
Setting Metric LSTM Transformer Ensemble
3 Class Accuracy .6944 .7189 .7271
(-6.9 %) (-0.6 %) (-5.9 %)
F1 .3849 .5146 .4956
(-20.6 %) (-1.8 %) (-11.9 %)
2 Class Accuracy .7899 .8254 .8242
(-7.0 %) (-4.1 %) (-5.7 %)
F1 .6144 .7742 .7429
(-20.9 %) (-1.7 %) (-8.4 %)
It becomes apparent that even our simple participant representation that we incorporate into our models contributes positively to individual alignment.
6 Experiments II - Perception-Guided Multimodal Transformer
@@
We propose a new way to combine content and sequential information by adapting a Transformer model. The novel idea is to directly apply the fixation sequences in form of transition matrices by adding them to the attention weights in the multimodal Transformer model. The intention behind this is to find an alternative way of combining the contextual and the sequential data in one model. We call this a Perception-Guided Multimodal Transformer (PGMT). An advantage of this procedure is that it does not require an increased number of trainable parameters when compared to the vanilla transformer approach. Table 5 shows the transition matrix derived from the first 5 rows of table 1.
Table 5: Transition Matrix (partial) for Participant EWCX and Stimulus ID 2412873.
vis_wall txt_wall off txt_rock
vis_wall 0 1 0 0
txt_wall 0 0 1 0
off 0 0 0 1
txt_rock 0 1 0 0
The attention manipulation process that we chose is straight-forward. We use the src_mask parameter in the PyTorch Transformer Encoder to add the transition matrix to the attention weights. [^7] Note that although the parameter is named src_mask, the documentation states that If a FloatTensor is provided, it will be added to the attention weight The attention weight refers to the term in brackets in equation 1 [24].
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
@@
Due to the sparsity of the matrices, and the final softmax application that happens after adding the matrices to the weights, it is beneficial to scale the weights to ensure that the perception information is not vanishing during calculations. This is enabled by the manual tuning parameter $\lambda$. We found it also beneficial to transpose the matrix along the diagonal and add the result to the original matrix. Equation 2 shows this amplification function. An example based on the matrix in table 5 is shown below.
@@
When we input the matrix from table 5 into equation 2 and set the weight variable $\lambda$ to 5, we get Equation 3:
@@
This operation reduces the sparsity of the matrix and in combination with the upscaling lead to the best results. When evaluated on the test data, the PGMT achieves an on-par performance with the Ensemble model in both settings. Table 6 shows that the approach we chose leads to competitive results, with the advantage of needing less trainable parameters.
Table 6: Comparison of the Perception-Guided Multimodal Transformer with the Ensemble model.
Setting Model Accuracy F1
3 Classes PGMT .7761 .5294
Ensemble .7728 .5628
2 Classes PGMT .8715 .8196
Ensemble .8742 .8112
In this setup, we used 128 feed-forward dimensions in the transformer, a batch size of 16 and a learning rate of 0.0001. In this setting, the transition matrices are weighted by a $\lambda$ of 500, which we empirically determined to be most effective in this scenario.
7 Experiments III - POV Learning with GPT4
With GPT4 being the state-of-the-art Mulitmodal Large Language Model we investigate how well it performs on POV tasks like PCE and if it is able to utilize perception information to improve individual alignment through few-shot learning.
In our GPT4 experiments, we investigate i) how well a foundation model like GPT4 is capable of solving our PCE task and ii) if few-shot incontext learning can improve its performance. As before, the task for the model is to predict how the individual participants evaluated the question ’Does the caption mention the central entities in the image?’, given a certain stimulus. As with the experiments before, there are three possible outcomes: Yes, No, Unclear. The only difference is, that we now can prompt the model directly, without optimization on this task. The different setups we investigated vary in the amount of participant-specific information that is provided in the prompt. The data and the training and test splits are the same as before. The different setups are as follows:
@@ 7 Experiments III - POV Learning with GPT4
No subject-specific information is provided. The model is tasked to evaluate the question based only on the stimulus, completely ignoring individual assessments of stimuli.
@@
In addition to the stimulus itself, the subject-specific fixation sequence for the current stimulus is provided in the prompt.
@@
This setup provides an additional training sample as an example of the current individual’s behavior and adds it to the system prompt. Thus, each prompt additionally contains a stimulus, a corresponding fixation sequence and the individual response.
Two-shot:
In this setup, an additional second training sample is provided for each prompt. The second sample was chosen to contain a different individual response than the first sample to provide both, a positive and a negative sample.
@@
As before, we report Accuracy and the F1 score in both the 2 class (Unclear class removed for evaluation) and the 3 class setting as performance metrics for the experiments. The experiment was realised with requests to the OpenAI API.
Table 7: Experimental Results for GPT4 alignment.
Setting Metric Zero-shot Fixations One-shot Two-shot
3 Class Accuracy .6568 .6732 .6584 .6683
F1 .3220 .3265 .3132 .3250
2 Class Accuracy .7472 .7657 .7490 .7630
F1 .5156 .5225 .5018 .5210
@@
The are two noteworthy outcomes regarding our GPT4 research questions: Concerning ii) it appears that GPT4 cannot benefit from few-shot learning for the PCE task. At least not in the way we provided fixation information and demonstrations of individual behavior in the prompt. The stimulus might be too complex, for only 1 or 2 training samples. Still, incontext learning with the fixation sequence only improves the performance. Concerning i) we found that overall GPT4’s performance in PCE is considerably and consistently lower than all approaches presented before. This suggests that GPT4 has not been instruction-tuned on a task similar to PCE and that it struggles with this task in general and with individual alignment in particular. Obviously, further research is needed to assess if post-training helps for POV tasks.
8 Conclusions
@@
In this paper we introduce POV Learning. By demonstrating the potential of taking a subject’s point-of-view and incorporating individual characteristics into machine learning models, we can improve predictive performance on an individual level. We also showed that state-of-the-art foundation models like GPT4 struggle with alignment on an individual level on our newly introduced Perception-guided Crossmodal Entailment (PCE) task.
@@
In three different experiments, we explored different ways of using human perception sequences as alignment signals for POV-learning tasks. To determine the relevance of the information contained in a fixation sequence, we implemented an LSTM model that processes a purely symbolic representation of the AOIs in the order of the fixations. This information alone leads to solid results in the PCE task, with an accuracy value of .7466, compared to .6715 for a naive model and .7238 for a content-only transformer model. Combining the LSTM with a Transformer that processes additional contextual information in form of pre-trained features of the text and image batches of the multimodal stimulus, the resulting Ensemble model achieves a further improvement up to an Accuracy of .7728. From this experiment we conclude that the usage of human perception data leads to improved individual alignment in the PCE task.
@@
Additionally, we propose the Perception-Guided Multimodal Transformer (PGMT), which manipulates the attention weights according to sequence information, which improves the Accuracy to .7761. In comparison to the Ensemble model, this approach has the advantage that the extra parameters of the LSTM model can be avoided while the number of parameters in the Transformer is the same.
With an Accuracy of only .6732, GPT4 is clearly outperformed by the PGMT and struggles with both, the PCE tasks and in-context learning in this setting.
In addition, we conduct an ablation study of the first experiments to determine how much the vector representations of the individual participants contribute to the performance. This ablation shows that even a rudimentary representation improves alignment by a considerable margin. Without the individual representations, the model performance decreases up to 20 percent. This further demonstrates the importance of individual alignment signals.
@@
In summary, we presented a new benchmark data set for the novel Perception-guided Crossmodal Entailment task PCE for POV Learning and proposed the PGMT model which can be used to explore the possibilities of aligning human-like perception with transformer models.

### v2

@@ 4.2 Perception-Guided LSTM Model
For this setup, we implement a unidirectional LSTM model that only receives symbolic representations of the AOIs as a sequence, called $x^{\text{\tiny ALIGN}_{\text{eye}}}$. In our case, we refer to all annotated regions of the eye tracking study as AOIs. If we take the sequence from Table 4 as an example, each row corresponds to one step in the LSTM. In this example, the symbol denoted vis_wall would correspond to the input in the first step t1, txt_wall to t2, and so on until the last element of the sequence vis_wall is reached. Before feeding the sequence into the LSTM, we put the one-hot encoded elements through an embedding layer that outputs a dense low-dimensional vector representation of the elements in the AOI sequence. Optionally, we also apply the same technique to create embeddings for the participants, called $x^{\text{\tiny ALIGN}_{\text{id}}}$. As a final step, we concatenate each AOI sequence element with the current participant embedding and put it through a feedforward layer to receive a combined representation, capturing Who is looking at which AOI. This final representation is then used as input for the LSTM, of which we take the final hidden state and perform the classification $\hat{y}^{\text{\tiny POV}}$ by running it through a feedforward layer. During training, the embedding layers are also updated, so that the resulting representations are fitted to the task. Overall, there are 837 unique AOI identifiers (i.e., the text representations like ”vis_wall” or ”txt_zebra”) in the dataset. This setting can be seen as using only the individual alignment signal $x^{\text{\tiny ALIGN}}$, without the stimulus content information beyond its symbolic abstraction $x^{\text{\tiny STIM}}$.
@@ 4.3 Content-based Transformer Model
The fundamental building block for this baseline is a transformer encoder, which comprises multiple attention heads and layers. Figure 6 shows the baseline model, which can be extended and adapted to different tasks by adding additional layers on top. We use pre-trained BERT [^2] https://huggingface.co/docs/transformers/v4.28.1/en/model_doc/bert [6] embeddings that are obtained from the Huggingface Transformers Package [27] as representations for the linguistic inputs (the captions) and a pre-trained ResNet50333https://pytorch.org/vision/main/models/resnet.html [28] to obtain image features of the predefined regions.
[…]
wallvis walltxt off rocktxt
wallvis 0 1 0 0
walltxt 0 0 1 0
off 0 0 0 1
rocktxt 0 1 0 0
Figure 5: Transition Matrix (partial) for Participant EWCX and Stimulus ID 2412873.
[…]
@@
For the experiments described in the following, we additionally concatenated a low-dimensional participant representation to the cross-modal classification token, as a rudimentary alignment signal. There is a single feed-forward layer between the ResNet image features and the transformer encoder, which is used to downsize the image feature dimension to match the BERT Embedding dimension. For the final decision $\hat{y}^{\text{\tiny POV}}$, the classification token is put through a single feed-forward layer with three output dimensions that correspond to the three possible outcomes. Note that the ResNet and BERT weights are not updated during training. Conceptually, the main advantage of the Transformer architecture is that it can use background knowledge in form of the pre-trained representations of the input modalities to solve the Perception-guided Crossmodal Entailment task.
@@ 4.4 Ensemble Model
In addition to the perception-guided approach of the LSTM and the content-based approach of the Transformer model, we introduce a meta-model as a combination of both. In this ensemble model, we simply combine both estimates of the models and use the average as our final classification $\hat{y}^{\text{\tiny POV}}$.
5 Experiments I - The Benefit of Alignment Signals in Basic POV Models
In a first set of experiments, we investigated how the different types of information (content $x^{\text{\tiny STIM}}$ vs. perception $x^{\text{\tiny ALIGN}})$ contribute to solving the PCE task when using the basic ML models introduced in Sect. 4.2 and Sect. 4.3. For a fair comparison, all three approaches are trained using the same hyperparameters, based on the best-performing ensemble model.
@@ 5 Experiments I - The Benefit of Alignment Signals in Basic POV Models
We use 6 attention heads and layers for the transformer, 256 dimensions for the transformer feed-forward network and 32 embedding dimensions (for both the participant and AOI embedding) in the LSTM network. All three models are trained using a batch size of 128 and a learning rate of 0.0001. We report the prediction accuracy and the F1 score (averaged over all classes) on the test split as evaluation metrics $g^{\text{\tiny POV}}$ for all conducted experiments. Due to the rarity of the unclear class, we also report accuracy and F1 scores that exclude samples with this class (’2 Class’ setting). In addition to our models, we report a naive baseline that only predicts the most frequent class. Table 1 shows the effect of different combinations of sequential and contextual information and functions as an ablation. It can be cleary seen that a combination of all available signals lead to the best results.
@@
These results clearly confirm our hypothesis that the fixation sequence contains information that can be used as an alignment signal $x^{\text{\tiny ALIGN}_{\text{eye}}}$ in a POV learning task. The naive baseline only reached half of the F1 score of the worst model. In addition, they show that the combination of $x^{\text{\tiny STIM}}$ and sequential information in an ensemble model leads to the best results, indicating that they complement each other.
To test the importance of providing a participant id $x^{\text{\tiny ALIGN}_{\text{id}}}$ to enable the personalization of the alignment signal, we investigate how much the individual participant embeddings (see Sect. 4.2) of our models contribute to the overall performance. It becomes apparent that even the simple representation of the participants that we learn within our models provides a useful alignment signal $x^{\text{\tiny ALIGN}_{\text{id}}}$ and improves the individual alignment.
Table 1: Effects of the presence of alignment signals
Model Eyetrack User Stimulus 3 classes 2 classes
$f()$ $x^{\text{\tiny ALIGN}_{\text{eye}}}$ $x^{\text{\tiny ALIGN}_{\text{id}}}$ $x^{\text{\tiny STIM}}$ F1 Acc. F1 Acc.
Naive .2678 .6715 .4330 .7639
LSTM + .3849 .6944 .6144 .7899
LSTM + + .4849 .7466 .7767 .8494
Transformer + .5146 .7189 .7742 .8254
Transformer + + .5240 .7238 .7879 .8611
Ensemble + + + .5628 .7728 .8112 .8742
6 Experiments II - The Perception-Guided Multimodal Transformer
@@
We propose a more efficient approach to combining content and sequential information by directly adding the fixation sequences in the form of transition matrices to the attention weights in the multimodal Transformer model. The basic idea behind this is to guide the attention heads to concentrate on AOI pairs that appear consecutively in the perception traces. We call this a Perception-Guided Multimodal Transformer (PGMT). An additional advantage of this procedure is that it requires a reduced number of trainable parameters when compared to the vanilla transformer approach and no additional model as the LSTM in the ensemble approach. Table 6 shows the transition matrix derived from the first 5 rows of table 4.
@@
We use the src_mask parameter in the PyTorch Transformer Encoder to add the transition matrix to the attention weights.Due to the sparsity of the matrices and the final softmax application that happens after adding the matrices to the weights, it is beneficial to scale the weights to ensure that the perception information is not vanishing during calculations. This is enabled by the manual tuning parameter $\lambda$. We also found it beneficial to transpose the matrix along the diagonal and add the result to the original matrix.
@@
An example based on the matrix in Table 6 is shown below: When we input the matrix from table 6 and set the weight variable $\lambda$ to 5, we get:
@@
This operation reduces the sparsity of the matrix and in combination with upscaling leads to the best results.
In this setup, we used 128 feed-forward dimensions in the transformer, a batch size of 16 and a learning rate of 0.0001. The transition matrices are weighted by $\lambda$ $=500$, which was manually tuned and was most effective in this scenario. When evaluated on the test data, the PGMT achieves an on-par performance with the Ensemble model in both settings. Table 8 shows that the approach leads to competitive results, while requiring less trainable parameters and no separate model to capture sequential information as in the ensemble.
7 Experiments III - In-Context POV Learning with LLMs
So far, we have shown that different ML models can be efficiently trained to benefit from POV signals $x^{\text{\tiny ALIGN}}$ resulting in improved predictive performance. Large Language Models (LLMs) have the potential to generate predictions, without training their parameters, just by prompting them with demonstrations of $x^{\text{\tiny ALIGN}}$.
With two state-of-the-art multi-modal large language models, we investigate how well they perform on POV tasks like PCE and if they are able to utilize perception information to improve individual alignment through few-shot learning (i.e., without any form of post-training). We investigate i) how well multimodal large language models like GPT4 and Pixtral are capable of solving our PCE task, and ii) if few-shot in-context learning can improve their performance. As before, the task for the model is to predict how the individual participants evaluated the question ’Does the caption mention the central entities in the image?’, given a certain stimulus. As in the experiments before, there are three possible outcomes: Yes, No, Unclear. The different setups vary in the amount of participant-specific information $X^{\text{\tiny ALIGN}}$ provided in the prompt. The data and the training and test splits are the same as before:
@@ 7 Experiments III - In-Context POV Learning with LLMs
No participant-specific information $x^{\text{\tiny ALIGN}}$ is provided. The model is tasked with evaluating the question based only on the stimulus $x^{\text{\tiny STIM}}$, completely ignoring individual evaluations of the stimuli.
@@
In addition to the stimulus itself, the subject-specific fixation sequence for the current stimulus $x^{\text{\tiny ALIGN}}$ is provided in the prompt.
@@
This setup provides an additional training sample $x^{\text{\tiny ALIGN}}$ as a demonstration of the current individual’s behavior and adds it to the system prompt. Thus, each prompt additionally contains a stimulus, a corresponding fixation sequence, and the individual response.
@@
As before, we report Accuracy and the F1 score in both the 2 class and the 3 class setting as performance metrics for the experiments. The experiment was carried out with requests to the OpenAI API and with a local instance of Pixtral [^4] https://huggingface.co/mistralai/Pixtral-12B-2409.
[…]
Setting Metric Result
Ensemble Acc. .8742
(2 Class) F1 .8112
PGMT Acc. .8715
(2 Class) F1 .8196
Ensemble Acc. .7728
(3 Class) F1 .5628
PGMT Acc. .7761
(3 Class) F1 .5294
Figure 7: Comparison of the Perception-Guided Multimodal Transformer with the Ensemble model.
[…]
Setting Metric Zero-shot Fixation One-shot
Pixtral Acc. .7340 .7439 .5180
(2 Class) F1 .5208 .4985 .4485
GPT4 Acc. .7472 .7657 .7490
(2 Class) F1 .5156 .5225 .5018
Pixtral Acc. .6454 .6535 .4460
(3 Class) F1 .3251 .3110 .2770
GPT4 Acc. .6568 .6732 .6584
(3 Class) F1 .3220 .3265 .3132
Figure 8: Experimental Results for LMM alignment with GPT-4 and Pixtral 12B.
@@
There are two noteworthy outcomes regarding our LMM research questions: Concerning ii) it appears that the LLMs cannot benefit from few-shot learning for the PCE task. At least not in the way we provided fixation information and demonstrations of individual behavior $X^{\text{\tiny ALIGN}}$ in the prompt. A perception trace and the PCE task might be too different and complex to tasks commonly used in post-training. However, in-context learning with the fixation sequence only improves performance. Concerning i) we found that overall LLM performance in PCE is considerably and consistently lower than all approaches presented before. This suggests that the tested LLMs have not been instruction-tuned on a task similar to PCE and that they struggle with this task in general and with individual alignment in particular. Obviously, more research is needed to assess if post-training helps for POV tasks.
8 Discussion
@@
In this paper, we introduce and formalize POV Learning. By demonstrating the potential of taking a subject’s point-of-view and incorporating individual characteristics into machine learning models, we can improve predictive performance on an individual level. We also showed that state-of-the-art multimodal models like GPT4 struggle with alignment on an individual level on our newly introduced Perception-guided Crossmodal Entailment (PCE) task.
As opposed to alignment techniques like RLHF [19], which tackles alignment as a population-level problem, our work interprets alignment on an individual level.
@@
In three experiments, we explored different ways of using human perception sequences and symbolic representations based on user IDs as alignment signals for POV-learning tasks. To determine the relevance of the information contained in a fixation sequence, we implemented an LSTM model that processes a purely symbolic representation of the AOIs in the order of the fixations. This information alone leads to solid results in the PCE task, with an accuracy value of .7466, compared to .6715 for a naive model and .7238 for a content-only transformer model. Combining the LSTM with a Transformer that processes additional contextual information in the form of pre-trained features of the text and image batches of the multimodal stimulus, the resulting Ensemble model achieves a further improvement up to an accuracy of .7728. From this experiment, we conclude that the usage of human perception data leads to improved individual alignment in the PCE task.
@@
Additionally, we propose the Perception-Guided Multimodal Transformer (PGMT), which manipulates the attention weights according to sequence information, which improves Accuracy to .7761. In comparison to the Ensemble model, this approach has the advantage that the extra parameters of the LSTM model can be avoided while the number of parameters in the Transformer stays the same. With a maximum Accuracy of .6732, the LLMs are clearly outperformed by the PGMT and struggle with both the PCE tasks and in-context learning in this setting.
@@
In summary, we provided a new benchmark data set for the novel Perception-guided Crossmodal Entailment task PCE for POV Learning and proposed the PGMT model, which can be used to explore the possibilities of aligning human-like perception with transformer models by exploiting individual alignment signals.